Shift report 20000818 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Aug 18 2000 - 23:58:39 EDT

  • Next message: http account: "Shift report 20000819 08:00-16:00"

    ------------------------------------------------------------
    Date: Fri Aug 18 23:58:39 2000, Shift: 16:00-24:00
    Supervisor: 
    EJK, with: DB, RD, mouse/rain
    Log book#4 pages 117 - 122
    
    Shift Summary:
    ------------------------------------------------------------
    1. Quench link interlock for a hour.
       Beam was cogged from 22:10 to 22.40, and then
       message was "Beam will be dumped : 
       preparing ring entry for cryo".
    
    2. 30 minitues beam, RUN 2348, about 5K collision.
    
    3. Ramiro swapped a FEE card at 17:10
    
       DB found scaler 5 counts which is trigger 1 at "NO BEAM".
       He cycled crate power and problem went away.(p119) 
    
       At 22:30, T1 trip, Ramiro came.
          - channel G16,17,18 failed to load HV.
          - All Voltage in T1 are OFF!!!  
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Aug 19 2000 - 00:02:12 EDT